### v3

# A note on generalized tensor CUR approximation for tensor pairs and tensor triplets based on the tubal product

Abstract
In this note, we briefly present a generalized tensor CUR (GTCUR) approximation for tensor pairs $(\underline{\bf X},\underline{\bf Y})$ and tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ based on the tubal product (t-product). We use the tensor Discrete Empirical Interpolation Method (TDEIM) to do these extensions. We show how the TDEIM can be utilized to generalize the classical tensor CUR (TCUR) approximation, which acts only on a single tensor, to jointly compute the TCUR of two and three tensors. This approach can be used to sample relevant lateral/horizontal slices of one data tensor relative to one or two other data tensors. For some special cases, the Generalized TCUR (GTCUR) method is reduced to the classical TCUR for both tensor pairs and tensor triplets in a similar fashion as shown for the matrices.
keywords:
CUR approximation, generalized tensor SVD, tubal product
MSC:
15A69 , 46N40 , 15A23
journal: Elsevier
[…]
1 Introduction
The matrix CUR (MCUR) approximation is a well-known technique for fast low-rank approximation of matrices [1]. It offers an interpretable approximation and a more compact data representation as it uses the actual columns/rows of the underlying data matrix to build the factor matrices. The generalized MCUR (GMCUR) approximation for matrix pairs and matrix triplets were presented in [2] and [3], respectively. Similar to the SVD, the MCUR is applied to a single matrix while the GMCUR considers a pair and a triplet of matrices and provides low-rank approximations of two/three matrices simultaneously. The key idea behind this generalization is using the same column indices for two given data matrices. To be more precise, let two data matrices ${\bf X}\in\mathbb{R}^{I_{1}\times I_{2}},\,(I_{1}\geq I_{2}),\,{\bf Y}\in\mathbb{R}^{I_{3}\times I_{2}},\,(I_{3}\geq I_{2})$ that have the same number of columns be given. The generalized SVD (GSVD) [4, 5], guarantees the following decompositions
${\bf X}$ $=$ ${\bf U}\,{\rm diag}(\alpha_{1},\cdots,\alpha_{I_{2}})\,{\bf Z},\quad\alpha_{i}\in[0,1]$ (1)
${\bf Y}$ $=$ ${\bf V}\,{\rm diag}(\beta_{1},\cdots,\beta_{I_{2}})\,{\bf Z},\quad\beta_{i}\in[0,1]$ (2)
where $\alpha_{i}^{2}+\beta_{i}^{2}=1$ with the ratios $\beta_{i}/\alpha_{i}$ of nondecreasing order for $i=1,2,\ldots,I_{2}$. Here, ${\bf U}\in\mathbb{R}^{I_{1}\times I_{1}},\,{\bf V}\in\mathbb{R}^{I_{3}\times I_{3}}$ are orthogonal, while ${\bf Z}\in\mathbb{R}^{I_{2}\times I_{2}}$ is a nonsingular matrix. As can be seen, the GSVD shares the common factor matrix ${\bf Z}$ between the given data matrices ${\bf X}$ and ${\bf Y}$. On the other hand, from the theory of the MCUR approximation, we know that by applying the Discrete Empirical Interpolation Method (DEIM) [6] to the basis ${\bf Z}$ (to be discussed in Section 3), we can sample appropriate column indices from the data matrices ${\bf X}$ and ${\bf Y}$. Since the same basis matrix ${\bf Z}$ is used within the DEIM, the same indices will be produced for sampling the columns of the matrix ${\bf X}$ and the matrix ${\bf Y}$. Besides, the matrices ${\bf U}$ and ${\bf V}$ can be used to sample row indices of the matrices ${\bf X}$ and ${\bf Y}$, respectively. This means that the selected row indices for these matrices can be different. As a result, for two matrices ${\bf X}$ and ${\bf Y}$ with the same number of columns, we can use the identical column indices for their MCUR approximations. The authors in [2] show that this idea can be potentially used in applications where one is interested in extracting the most discriminative features from one data set relative to another data set. It is also motivated for recovering data matrices perturbed with colored noise and subgroup discovery application. Following this idea, in [3], the authors consider triplet tensors $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ and use the restricted SVD [7] to build their common factor matrices. Then, the DEIM is applied to the common factor matrices to sample columns and rows. It is shown that this approach can be used to extract the most discriminative features from one data set relative to another data set. The GMCUR for tensor triplets is also motivated by the canonical correlation analysis and was successfully used for recovering datasets perturbed with colored noise, See [2, 2] for details.
In this note, due to the applications of tensors [8, 9, 10], we focus on the tensor SVD (t-SVD) [11] that is defined based on the tubal product (t-product). The t-SVD has similar properties as the classical SVD. In particular, in contrast to the Tucker decomposition [12, 13] or Canonical polyadic decomposition [14, 15], its truncation provides the best low tubal rank approximation in the least-squares sense. The GSVD was generalized to tensors based on the t-SVD in [16, 17]. Motivated by the works [2, 3], we extend the tensor CUR (TCUR) approximation [18] to a pair and a triplet of tensors based on the t-product and we refer to as the generalized TCUR (GTCUR). As real-world datasets often have multidimensional structure, it is more favorable to generalize the ideas proposed in [2, 19] to tensors so that the structure of the datatensors are preserved. Indeed, we aim at achieving this target and present several generalizations of GMCUR to tensors. To this end, we use the tensor DEIM (TDEIM) index selection algorithm [20], which we recently developed to sample important lateral and horizontal slices, and this plays a key role in the development of the GTCUR method. Similar to the matrix case, the TDEIM helps to compute the same lateral/horizontal slices for a given pair or triplet of third order tensors. We remark that alternative sampling procedures, such as top tubal leverage scores [21, 18] or its sampling version, may be employed in place of the TDEIM, but because the TDEIM has been shown to give the optimal sample method [6], we adopt it in our research.
The remainder of this note is organized as follows. We first present some basic tensor concepts in Section 2. The t-SVD is introduced in Section 2.1 with some intuitions behind this model. The matrix cross approximation is outlined in Section 3 and the Tensor Discrete Empirical Interpolation Method (TDEIM) as an extension of the DEIM method is presented in Section 4. In Sections 5 and 6, we show how the matrix cross approximation for matrix pairs/triplets can be generalized to tensor pairs and tensor triplets, respectively. The conclusion is given in Section 7.
2 Preliminaries
Before presenting the main materials, we first introduce the basic notations and definitions. An underlined bold capital case letter, a bold capital case letter, and a bold lower case letter are used to represent a tensor, a matrix, and a vector, respectively. For a third-order tensor $\underline{{\bf X}},$ we call the slices $\underline{{\bf X}}(:,:,k),\,\underline{{\bf X}}(:,j,:),\,\underline{{\bf X}}(i,:,:)$ frontal, lateral, and horizontal slices. For simplicity, sometimes for a frontal slice $\underline{\bf X}(:,:,i)$, we use the notation $\underline{\bf X}_{i}$. For a third-order tensor $\underline{\bf X},$ the fiber $\underline{{\bf X}}(i,j,:)$ is called a tube. The matrix’s component-wise complex conjugate is denoted by the symbol “conj”. To indicate a subset of a matrix or tensor, we use the MATLAB notations. For instance, for a given data matrix ${\bf X}$, by ${\bf X}(:,\mathcal{J})$ and ${\bf X}(\mathcal{I},:)$ we mean two matrices sampling a part of the rows and columns of the matrix ${\bf X}$, respectively, where $\mathcal{I}\subset\{1,2,\ldots,I_{2}\}$ and $\mathcal{J}\subset\{1,2,\ldots,I_{1}\}$. We need to present Definitions 1-5 to introduce the tensor SVD (t-SVD) model.
Definition 1.
(t-product) Let $\underline{\mathbf{X}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\mathbf{Y}}\in\mathbb{R}^{I_{2}\times I_{4}\times I_{3}}$, the t-product $\underline{\mathbf{X}}*\underline{\mathbf{Y}}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}}$ is defined as follows
$\underline{\mathbf{C}}=\underline{\mathbf{X}}*\underline{\mathbf{Y}}={\rm fold}\left({{\rm circ}\left(\underline{\mathbf{X}}\right){\rm unfold}\left(\underline{\mathbf{Y}}\right)}\right),$ (3)
where
${\rm circ}\left(\underline{\mathbf{X}}\right)=\begin{bmatrix}\underline{\mathbf{X}}(:,:,1)&\underline{\mathbf{X}}(:,:,I_{3})&\cdots&\underline{\mathbf{X}}(:,:,2)\\ \underline{\mathbf{X}}(:,:,2)&\underline{\mathbf{X}}(:,:,1)&\cdots&\underline{\mathbf{X}}(:,:,3)\\ \vdots&\vdots&\ddots&\vdots\\ \underline{\mathbf{X}}(:,:,I_{3})&\underline{\mathbf{X}}(:,:,I_{3}-1)&\cdots&\underline{\mathbf{X}}(:,:,1)\end{bmatrix},$
and
${\rm unfold}(\underline{\mathbf{Y}})=\begin{bmatrix}\underline{\mathbf{Y}}(:,:,1)\\ \underline{\mathbf{Y}}(:,:,2)\\ \vdots\\ \underline{\mathbf{Y}}(:,:,I_{3})\end{bmatrix},\hskip 14.22636pt\underline{\mathbf{Y}}={\rm fold}\left({\rm unfold}\left(\underline{\mathbf{Y}}\right)\right).$
The Discreet Fourier Transform (DFT) is used to execute the t-product as described in [22, 11], and it was recommended in [23] to use any invertible transformation instead of the DFT. Later, in [24] and [25], nonivertible and even nonlinear mapping were employed. The ability to compute the t-SVD of a data tensor with a lower tubal rank is a benefit of using such unitary transformations [26, 24]. The MATLAB command ${\rm fft}(\underline{\bf X},[],3)$, computes the DFT of every tube of the data tensor $\underline{\bf X}$. The fast version of the t-product in which the DFT of only the first $\lceil\frac{I_{3}+1}{2}\rceil$ frontal slices is needed is summarized in Algorithm 2, see [27, 28] for details.
Definition 2.
(Transpose) The transpose of a tensor $\underline{\mathbf{X}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ is denoted by $\underline{\mathbf{X}}^{T}\in\mathbb{R}^{I_{2}\times I_{1}\times I_{3}}$ produced by applying the transpose to all frontal slices of the tensor $\underline{\mathbf{X}}$ and reversing the order of the transposed frontal slices from the second until the last one.
Definition 3.
(Identity tensor) Identity tensor $\underline{\mathbf{I}}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}}$ is a tensor whose first frontal slice is an identity matrix of size $I_{1}\times I_{1}$ and all other frontal slices are zero. It is easy to show $\underline{{\bf I}}*\underline{{\bf X}}=\underline{{\bf X}}$ and $\underline{{\bf X}}*\underline{{\bf I}}=\underline{{\bf X}}$ for all tensors of conforming sizes.
Definition 4.
(Orthogonal tensor) A tensor $\underline{\mathbf{X}}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}}$ is orthogonal if ${\underline{\mathbf{X}}^{T}}*\underline{\mathbf{X}}=\underline{\mathbf{X}}*{\underline{\mathbf{X}}^{T}}=\underline{\mathbf{I}}$.
Definition 5.
(f-diagonal tensor) If all frontal slices of a tensor are diagonal, then the tensor is called f-diagonal.
Definition 6.
(Moore–Penrose pseudoinverse of a tensor) Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ be given. The Moore-Penrose (MP) pseudoinverse of the tensor $\underline{\bf X}$ is denoted by $\underline{\bf X}^{{\dagger}}\in\mathbb{R}^{I_{2}\times I_{1}\times I_{3}}$ is a unique tensor satisfying the following four equations:
$\underline{\bf X}^{{\dagger}}*\underline{\bf X}*\underline{\bf X}^{{\dagger}}=\underline{\bf X}^{{\dagger}},\quad\underline{\bf X}*\underline{\bf X}^{{\dagger}}*\underline{\bf X}=\underline{\bf X},$
$(\underline{\bf X}*\underline{\bf X}^{{\dagger}})^{T}=\underline{\bf X}*\underline{\bf X}^{{\dagger}},\quad(\underline{\bf X}^{{\dagger}}*\underline{\bf X})^{T}=\underline{\bf X}^{{\dagger}}*\underline{\bf X}.$
The MP pseudoinverse of a tensor can also be computed in the Fourier domain and this is shown in Algorithm 2.
Input : Two data tensors $\underline{\mathbf{X}}\in{\mathbb{R}^{{I_{1}}\times{I_{2}}\times{I_{3}}}},\,\,\underline{\mathbf{Y}}\in{\mathbb{R}^{{I_{2}}\times{I_{4}}\times{I_{3}}}}$
Output : t-product $\underline{\mathbf{C}}=\underline{\mathbf{X}}*\underline{\mathbf{Y}}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}}$
1 $\widehat{\underline{\mathbf{X}}}={\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$;
2 $\widehat{\underline{\mathbf{Y}}}={\rm fft}\left({\underline{\mathbf{Y}},[],3}\right)$;
3 for $i=1,2,\ldots,\lceil\frac{I_{3}+1}{2}\rceil$ do
4       $\widehat{\underline{\mathbf{C}}}\left({:,:,i}\right)=\widehat{\underline{\mathbf{X}}}\left({:,:,i}\right)\,\widehat{\underline{\mathbf{Y}}}\left({:,:,i}\right)$;
5
6 end for
7for $i=\lceil\frac{I_{3}+1}{2}\rceil+1\ldots,I_{3}$ do
8       $\widehat{\underline{\mathbf{C}}}\left({:,:,i}\right)={\rm conj}(\widehat{\underline{\mathbf{C}}}\left({:,:,I_{3}-i+2}\right))$;
9 end for
$\underline{\mathbf{C}}={\rm ifft}\left({\widehat{\underline{\mathbf{C}}},[],3}\right)$;
Algorithm 1 Fast t-product of two tensors [22, 28]
Input : The data tensor $\underline{\mathbf{X}}\in{\mathbb{R}^{{I_{1}}\times{I_{2}}\times{I_{3}}}}$
Output : Moore-Penrose pseudoinverse $\underline{\mathbf{X}}^{{\dagger}}\in\mathbb{R}^{I_{2}\times I_{1}\times I_{3}}$
1 $\widehat{\underline{\mathbf{X}}}={\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$;
2 for $i=1,2,\ldots,\lceil\frac{I_{3}+1}{2}\rceil$ do
3       $\widehat{\underline{\mathbf{C}}}\left({:,:,i}\right)={\rm pinv}\,(\widehat{{\bf X}}(:,:,i))$;
4
5 end for
6for $i=\lceil\frac{I_{3}+1}{2}\rceil+1,\ldots,I_{3}$ do
7       $\widehat{\underline{\mathbf{C}}}\left({:,:,i}\right)={\rm conj}(\widehat{\underline{\mathbf{C}}}\left({:,:,I_{3}-i+2}\right))$;
8 end for
$\underline{\mathbf{X}}^{{\dagger}}={\rm ifft}\left({\widehat{\underline{\mathbf{C}}},[],3}\right)$;
Algorithm 2 Fast Moore-Penrose pseudoinverse computation of the tensor $\underline{\bf X}$
2.1 Tensor decompositions based on the t-product and tubal leverage-scores
The tensor SVD (t-SVD) is a viable tensor decomposition that represents a tensor as the t-product of three tensors. The first and last tensors are orthogonal, while the middle tensor is an f-diagonal tensor. Let $\underline{{\bf X}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, then the t-SVD gives the following model:
$\underline{{\bf X}}\approx\underline{\bf U}*\underline{\bf S}*\underline{\bf V}^{T},$
where $\underline{\bf U}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf S}\in\mathbb{R}^{R\times R\times I_{3}},$ and $\underline{\bf V}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$. The tensors $\underline{\bf U}$ and $\underline{\bf V}$ are orthogonal, while the tensor $\underline{\bf S}$ is f-diagonal. The procedure of the computation of the t-SVD is presented in Algorithm 3.
Input : The data tensor $\underline{\mathbf{X}}\in{\mathbb{R}^{{I_{1}}\times{I_{2}}\times{I_{3}}}}$ and a target tubal-rank $R$
Output : The truncated t-SVD of the tensor $\underline{\bf X}$ as $\underline{\bf X}\approx\underline{\bf U}*\underline{\bf S}*\underline{\bf V}^{T}$
1 $\widehat{\underline{\mathbf{X}}}={\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$;
2 for $i=1,2,\ldots,\lceil\frac{I_{3}+1}{2}\rceil$ do
3       $[\widehat{\underline{\mathbf{U}}}\left({:,:,i}\right),\widehat{\underline{\bf S}}(:,:,i),\widehat{{\bf V}}(:,:,i)]={\rm Truncated\operatorname{-}svd}\,(\widehat{{\bf X}}(:,:,i),R)$;
4
5 end for
6for $i=\lceil\frac{I_{3}+1}{2}\rceil+1,\ldots,I_{3}$ do
7       $\widehat{\underline{\mathbf{U}}}\left({:,:,i}\right)={\rm conj}(\widehat{\underline{\mathbf{U}}}\left({:,:,I_{3}-i+2}\right))$;
8       $\widehat{\underline{\mathbf{S}}}\left({:,:,i}\right)=\widehat{\underline{\mathbf{S}}}\left({:,:,I_{3}-i+2}\right)$;
9       $\widehat{\underline{\mathbf{V}}}\left({:,:,i}\right)={\rm conj}(\widehat{\underline{\mathbf{V}}}\left({:,:,I_{3}-i+2}\right))$;
10 end for
$\underline{\mathbf{U}}_{R}={\rm ifft}\left({\widehat{\underline{\mathbf{U}}},[],3}\right)$; $\underline{\mathbf{S}}_{R}={\rm ifft}\left({\widehat{\underline{\mathbf{S}}},[],3}\right)$; $\underline{\mathbf{V}}_{R}={\rm ifft}\left({\widehat{\underline{\mathbf{V}}},[],3}\right)$;
Algorithm 3 The truncated t-SVD decomposition of the tensor
3 Matrix CUR (MCUR) approximation of a single matrix and matrix pairs
The Matrix CUR (MCUR), or cross approximation, is a well-known method for fast low-rank matrix approximation with interpretable factor matrices and linear computational complexity [1]. This technique proposes to choose the individual columns and rows of a matrix in order to preserve the original data matrix’s structure, such as negativity or sparsity. Let us describe it formally. Given ${\bf X}\in\mathbb{R}^{I_{1}\times I_{2}}$ be a give data matrix. The MCUR approximation seeks the approximation of the form ${\bf X}\approx{\bf C}{\bf U}{\bf R}$ where ${\bf C}={\bf X}(:,\mathcal{J}),\,{\bf R}={\bf X}(\mathcal{I},:)$ and $\mathcal{I}\subset\{1,2,\ldots,I_{1}\}$ and $\mathcal{J}\subset\{1,2,\ldots,I_{2}\}$. It is obvious that the best middle matrix is
${\bf U}={\bf C}^{{\dagger}}{\bf X}{{\bf R}}^{{\dagger}},$ (4)
as it minimizes the error term or
$\min_{{\bf U}}\|{\bf X}-{\bf C}{\bf U}{\bf R}\|^{2}_{F}.$ (5)
A deterministic or random selection of the columns and rows is possible, with the option of obtaining additive or relative approximation errors. It is generally known that the columns or rows with the highest volume might provide virtually optimal solutions in a deterministic scenario [29]. The DEIM is another deterministic way to select a matrix’s columns or rows and relies on the top singular vectors [6]. The three probability distributions that are most typically used to choose columns and rows are uniform, length-squared, and leverage-score distributions. For additional details on these sampling techniques, see [21]. It has been shown that approximations with relative error accuracy that are more practical in practice can be produced by sampling columns with the leverage-score probability distribution [30].
The MCUR approximation of matrix pairs (matrices with the same number of columns) as the extension of the classical MCUR approximations were presented in [2]. The key idea of this generalization is to select columns from two matrices that share the same column indices in order to extract the features of one dataset that are most pertinent to the other. Here, the common factor matrix produced by the GSVD of the two matrices is used to sample the columns and rows using the Discrete Empirical Interpolation Method (DEIM) technique. The MCUR approximation of triplet matrices was also presented in [3] , with the goal of identifying pertinent characteristics of one data matrix in relation to two other matrices. The authors in this instance apply the RSVD to produce two common factor matrices, to which the DEIM is then applied to sample the column and row indices. We first discuss the DEIM extension to tensors based on the t-product in the next section before using it to extend the TCUR technique to tensor pairs and tensor triplets in Sections 5 and 6, respectively.
4 Tensor Discrete Empirical Interpolation Method (TDEIM) for lateral/horizontal slice sampling
We have recently generalized the DEIM method to the tensor case based on the t-product [20] that we call tensor DEIM (TDEIM). It has been experimentally shown that the TDEIM achieved the best sampling accuracy compared to the existing sampling techniques, such as top tubal leverage score sampling and uniform sampling. The TDEIM to select horizontal slices is presented in Algorithm 4. A key concept in the TDEIM is the interpolatory projector, as we now define it. For a given set of indices ${\bf s}\in\mathbb{R}^{R}$, the full tubal-rank tensor [^1] A tensor with linear independent lateral slices, for example, the tensor $\underline{\bf U}$ obtained from the t-SVD can be used. $\underline{\bf U}\in\mathbb{R}^{I_{1}\times R\times I_{3}}$ and considering $\underline{\bf S}=\underline{\bf I}(:,{\bf s},:)\in\mathbb{R}^{I_{1}\times R\times I_{3}}$, as the selection tensor where $\underline{\bf I}\in\mathbb{R}^{R\times R\times I_{3}}$ is an identity tensor, we build the tensorial oblique projection operator defined as follows
$\underline{\mathcal{P}}=\underline{\bf U}*(\underline{\bf{S}}^{T}*\underline{\bf{U}})^{-1}*\underline{\bf{S}}^{T}.$ (6)
Given an arbitrary tensor $\underline{\bf G}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf X}=\underline{\mathcal{P}}*\underline{\bf G}$, we have
$\underline{\bf X}({\bf s},:,:)$ $=$ $\underline{\bf S}^{T}*\underline{\bf X}=\underline{\bf S}^{T}*\underline{\bf U}*(\underline{\bf{S}}^{T}*\underline{\bf{U}})^{-1}*\underline{\bf{S}}^{T}*\underline{\bf G}$ (7)
$=$ $\underline{\bf{S}}^{T}*\underline{\bf G}=\underline{\bf G}({\bf s},:,:).$
This means that the projection operator $\underline{\mathcal{P}}$ preserves the horizontal slices of $\underline{\bf G}$ specified by the index set ${\bf s}$. This justifies the name of interpolation. The TDEIM starts with selecting an index with the maximum Euclidean norm of the tubes in the first lateral slice of the basis, i.e. $\underline{\bf U}(:,1,:),$ to be the index of the first sampled horizontal slice. The subsequent indices are selected according to the indices with the maximum Euclidean norm of the tubes of the residual lateral slice that is computed by removing the direction of the tensorial interpolatory projection in the previous basis vectors from the subsequent one. To be more precise, let the indices ${\bf s}_{j-1}=\{s_{1},s_{2},\ldots,s_{j-1}\}$ have been already selected, and we want to select $s_{j}$. To do so, we compute the residual slice
$\underline{\bf R}(:,j,:)=\underline{\bf U}(:,j,:)-\underline{\mathcal{P}}_{j-1}*\underline{\bf U}(:,j,:),$
where $\underline{\mathcal{P}}_{j-1}=\underline{\bf U}^{j-1}*({\underline{\bf S}^{j-1}}^{T}*\underline{{\bf U}}^{j-1})^{-1}*{\underline{\bf S}^{j-1}}^{T},\,\underline{\bf S}^{j-1}=\underline{\bf I}(:,{\bf s}_{j-1},:),$ and $\underline{\bf U}^{j-1}=\underline{\bf U}(:,{\bf s}_{j-1},:)$. Then, a new index with the maximum Euclidean norm of tubes or
$s_{j}=\arg\max_{1\leq i\leq I_{1}}\|\underline{\bf R}(i,j,:)\|,$
is set as our new index $s_{j}$. This will be the new sampled horizontal slice’s index. The same procedure can be used to select lateral slices where the tensor $\underline{\bf V}$ obtained from the t-SVD of the tensor $\underline{\bf X}$ should be used.
Similar to the matrix case, let us introduce two quantities as follows
$\tilde{\eta}_{p}=\frac{1}{I_{3}}\max_{i}\,\left({\|(\widehat{\underline{\bf S}}^{T}_{i}\widehat{\underline{\bf U}}_{i})^{-1}\|_{2}^{2}}\right),\quad\tilde{\eta}_{q}=\frac{1}{I_{3}}\max_{i}\,\left({\|(\widehat{\underline{\bf V}}^{T}_{i}\widehat{\underline{\bf Q}}_{i})^{-1}\|_{2}^{2}}\right),$ (8)
that will be used in the next theorem. Here, $\underline{\bf V}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ is a bais tensor for the subspace of horizontal slices of the original data tensor, $\underline{\bf Q}=\underline{\bf I}(:,{\bf s}_{j-1},:)$ is a tensor of some sampled lateral slices of the identity tensor specified by the index set ${\bf s}_{j-1}=\{s_{1},s_{2},\ldots,s_{j-1}\},\,\widehat{\underline{\bf Q}}={\rm fft}(\underline{\bf Q},[],3),\,\widehat{\underline{\bf V}}={\rm fft}(\underline{\bf V},[],3),\,\widehat{\underline{\bf Q}}_{i}=\widehat{\underline{\bf Q}}(:,:,i),$ and $\widehat{\underline{\bf V}}_{i}=\widehat{\underline{\bf V}}(:,:,i)$.
The next theorem provides the upper error bound of the TCUR approximation obtained by the TDEIM sampling method.
Theorem 1.
[20] Suppose $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $1\leq R<\min(I_{1},I_{2})$. Assume that horizontal slice and lateral slice indices ${\bf p}$ and ${\bf q}$ give full tubal rank tensors $\underline{\bf C}=\underline{\bf X}(:,{\bf q},:)=\underline{\bf X}*\underline{\mathcal{Q}}$ and $\underline{\bf R}=\underline{\bf X}({\bf p},:,:)={\underline{\mathcal{P}}}*\underline{\bf X}$, where $\underline{\mathcal{P}}$ and $\underline{\mathcal{Q}}$ are tensorial interpolation projectors for horizontal and lateral slice sampling [^2] The tensorial interpolation projector for the lateral slice sampling is defined as $\underline{\mathcal{Q}}=\underline{{\bf V}}*(\underline{{\bf Q}}^{T}*\underline{{\bf V}})^{-1}*\underline{{\bf Q}}^{T}$, where $\underline{{\bf V}}$ is a basis tensor and $\underline{{\bf Q}}$ is the collection tensor., respectively with corresponding finite error constants $\tilde{\eta}_{p},\,\tilde{\eta}_{q}$ defined in (8) and set $\underline{\bf U}=\underline{{\bf C}}^{{\dagger}}*\underline{\bf X}*\underline{\bf R}^{{\dagger}}$. Then
$\|\underline{\bf X}-\underline{\bf C}*\underline{\bf U}*\underline{\bf R}\|^{2}_{F}\leq(\tilde{\eta}_{p}+\tilde{\eta}_{q})\sum_{i=1}^{I_{3}}\sum_{t>R}(\sigma^{i}_{t})^{2},$ (9)
where $\sigma^{i}_{t}$ is the $t$-th largest singular values of the frontal slice $\widehat{\underline{\bf X}}_{i}=\widehat{\underline{\bf X}}(:,:,i)$.
Theorem 1 shows that the TCUR approximation with the middle tensor defined above can provide approximation within a factor $\tilde{\eta}_{p}+\tilde{\eta}_{q}$ of the best tubal rank $k$ approximation. It also indicates that the conditioning of the problem depends on these two quantities and the lateral/horizontal slices should be selected in such a way that these quantities should be controlled.
Remark 2.
The basis tensors $\underline{\bf U}$ and $\underline{\bf V}$ required in Algorithm 4 can be computed very fast through the randomized truncated t-SVD [31, 32, 33]. This version can be regarded as a randomized version of the TDEIM algorithm.
Input :  $\underline{\bf U}\in\mathbb{R}^{I_{1}\times R\times I_{3}}$ with $R\leq I_{1}$ (linearly independent lateral slices)
Output : Indices ${\bf s}\in\mathbb{N}^{R}$ with distinct entries in $\{1,2,\ldots,I_{1}\}$
1 $s_{1}=\arg\max_{1\leq i\leq I_{1}}\|\underline{\bf U}(i,1,:)\|_{2}$
2 for $j=2,3,\ldots,R$ do
3       $\underline{{\bf C}}=\underline{\bf U}({\bf s},1:j-1,:)^{-1}*\underline{\bf U}({\bf s},j,:)$;
4       $\underline{{\bf R}}=\underline{\bf U}(:,j,:)-\underline{\bf U}(:,1:j-1,:)*\underline{{\bf C}}$;
5       $s_{j}=\arg\max_{1\leq i\leq I_{1}}\|\underline{{\bf R}}(i,j,:)\|_{2}$
6 end for
Algorithm 4 Tubal DEIM (TDEIM) index selection approach for horizontal slice selection
5 Tensor CUR (TCUR) approximation for tensor pairs based on the t-product
Inspired by the success of the MCUR approximation, several generalizations of these approaches to tensors have been proposed. For example, see [34] for the Tucker model, see [35, 36] for the TT decomposition, see [18] for the t-SVD, and see [37] for the TC/TR decomposition. In this note, we focus on the t-SVD case and extend it to a pair of tensors. The intuition behind this generalization is similar to the matrix case. More precisely, the indices of selected lateral slices are the same for two data tensors (see Figure 1 for a graphical illustration). The formal definition of the GTCUR for tensor pairs is presented below.
Figure 1: Visualization of the generalized tensor CUR (GTCUR) approximation for tensor pairs $(\underline{\bf X},\underline{\bf Y})$. The indices of the selected lateral slices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$ are identical while this is not necessarily true for the horizontal slices.
Definition 7.
Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$ be of full tubal rank, with $I_{1}\geq I_{2}$ and $I_{4}\geq I_{2}$. A GTCUR decomposition of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ with the tubal rank $R,$ is defined as follows
$\underline{\bf X}\approx{\underline{\bf C}}_{1}*\underline{\bf U}_{1}*\underline{\bf R}_{1}=(\underline{\bf X}*\underline{\bf P})*\underline{\bf U}_{1}*(\underline{\bf S}_{1}^{T}*\underline{\bf X}),$ (10)
$\underline{\bf Y}\approx{\underline{\bf C}}_{2}*\underline{\bf U}_{2}*\underline{\bf R}_{2}=(\underline{\bf Y}*\underline{\bf P})*\underline{\bf U}_{2}*(\underline{\bf S}_{2}^{T}*\underline{\bf Y}),$ (11)
where $\underline{\bf S}_{1}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf S}_{2}\in\mathbb{R}^{I_{4}\times R\times I_{3}}$ and $\underline{\bf P}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ are index the selection tensors ($R<I_{2}$).
Assume that ${\bf s}_{1},$ and ${\bf s}_{2}$ are the horizontal slice indices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$, respectively. Also, let ${\bf p}$ be the common lateral slice indices for the both tensors $\underline{\bf X}$ and $\underline{\bf Y}$. Then, formulations (10)-(11) can be equavalently written as
$\underline{\bf X}$$\approx$ $\underline{\bf X}(:,{\bf p},:)*\underline{\bf U}_{1}*\underline{\bf X}({\bf s}_{1},:,:),$ (12)
$\underline{\bf Y}$$\approx$ $\underline{\bf Y}(:,{\bf p},:)*\underline{\bf U}_{2}*\underline{\bf Y}({\bf s}_{2},:,:).$ (13)
The GTCUR for tensor pairs indeed computes the TCUR approximations of two given data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ and is motivated by the generalized t-SVD (GTSVD). For given data tensors $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$, the GTSVD guarantees the following decompositions [17]
$\underline{\bf X}$ $=$ $\underline{\bf U}*\underline{\bf C}*\underline{\bf Z}^{T},$ (14)
$\underline{\bf Y}$ $=$ $\underline{\bf V}*\underline{\bf S}*\underline{\bf Z}^{T},$ (15)
where $\underline{\bf U}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}},\,\underline{\bf V}\in\mathbb{R}^{I_{4}\times I_{4}\times I_{3}},\,\underline{\bf C}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}},\,\underline{\bf S}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}},\,\underline{\bf Z}\in\mathbb{R}^{I_{2}\times I_{2}\times I_{3}}$. Note that the tensors $\underline{\bf C}$ and $\underline{\bf S}$ are f-diagonal, and the tensors $\underline{\bf U}$ and $\underline{\bf V}$ are orthogonal, while the tensor $\underline{\bf Z}$ is nonsingular. The procedure of the computation of the GTSVD is presented in Algorithm 5. We need to apply the classical GSVD (lines 3-5) to the first $\lceil\frac{I_{3}+1}{2}\rceil$ frontal slices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$ in the Fourier domain and the rest of the slices are computed easily (Lines 6-12).
Input : The data tensors $\underline{\mathbf{X}}\in{\mathbb{R}^{{I_{1}}\times{I_{2}}\times{I_{3}}}}$ and $\underline{\mathbf{Y}}\in{\mathbb{R}^{{I_{4}}\times{I_{2}}\times{I_{3}}}}$
Output : The generalized t-SVD of $\underline{\mathbf{X}}$ and $\underline{\bf Y}$ as $\underline{\mathbf{X}}=\underline{\mathbf{U}}*\underline{\mathbf{C}}*\underline{\mathbf{Z}}$ and $\underline{\mathbf{Y}}=\underline{\mathbf{V}}*\underline{\mathbf{S}}*\underline{\mathbf{Z}}$
1 $\widehat{\underline{\mathbf{X}}}={\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$;
2 $\widehat{\underline{\mathbf{Y}}}={\rm fft}\left({\underline{\mathbf{Y}},[],3}\right)$;
3 for $i=1,2,\ldots,\lceil\frac{I_{3}+1}{2}\rceil$ do
4       $[\widehat{\underline{\mathbf{U}}}_{i},\,\widehat{\underline{\mathbf{V}}}_{i},\,\widehat{\underline{\mathbf{Z}}}_{i},\,\widehat{\underline{\mathbf{C}}}_{i},\,\widehat{\underline{\mathbf{S}}}_{i}]={\rm GSVD}\,(\widehat{\underline{{\bf X}}}(:,:,i),\widehat{\underline{{\bf Y}}}(:,:,i))$;
5
6 end for
7for $i=\lceil\frac{I_{3}+1}{2}\rceil+1,\ldots,I_{3}$ do
8       $\widehat{\underline{\mathbf{U}}}_{i}={\rm conj}(\widehat{\underline{\mathbf{U}}}_{I_{3}-i+2})$;
9       $\widehat{\underline{\mathbf{V}}}_{i}={\rm conj}({\underline{\mathbf{V}}}_{I_{3}-i+2})$;
10       $\widehat{\underline{\mathbf{Z}}}_{i}={\rm conj}(\widehat{\underline{\mathbf{Z}}}_{I_{3}-i+2})$;
11       $\widehat{\underline{\mathbf{C}}}_{i}=\widehat{\underline{\mathbf{C}}}_{I_{3}-i+2}$;
12       $\widehat{\underline{\mathbf{S}}}_{i}=\widehat{\underline{\mathbf{S}}}_{I_{3}-i+2}$;
13 end for
${\underline{\mathbf{U}}}={\rm ifft}\left({\widehat{\underline{\mathbf{U}}},[],3}\right)$; $\,{\underline{\mathbf{V}}}={\rm ifft}\left({\widehat{\underline{\mathbf{V}}},[],3}\right)$; $\,{\underline{\mathbf{Z}}}={\rm ifft}\left({\widehat{\underline{\mathbf{Z}}},[],3}\right)$; $\,{\underline{\mathbf{C}}}={\rm ifft}\left({\widehat{\underline{\mathbf{C}}},[],3}\right)$; $\,{\underline{\mathbf{S}}}={\rm ifft}\left({\widehat{\underline{\mathbf{S}}},[],3}\right)$;
Algorithm 5 Generalized t-SVD of $\underline{\bf X}$ and $\underline{\bf Y}$
We see that the GTSVD provides the same right tensor $\underline{\bf Z}$ in (14)-(15) and we can use it to sample lateral slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ based on the TDEIM algorithm. As a result, the same indices can be used to sample the lateral slices for the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$. The horizontal slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ can also be sampled using the left tensor parts $\underline{\bf U}$ and $\underline{\bf V}$, although they do not necessarily provide identical horizontal slice indices. Following this idea, we can compute the GTSVD of the tensors $\underline{\bf X},\,\underline{\bf Y}$ and by applying the TDEIM to the shared tensor factors $\underline{\bf U},\,\underline{\bf V},\,\underline{\bf Z}$, we can select indices of horizontal and lateral slices of the given data tensors. This approach is summarized in Algorithm 6. In Line 1 of Algorithm 6, we need to compute the GTSVD of two tensors and this is clearly prohibitive for large-scale tensors. However, to tackle this problem, the randomized algorithms proposed in [38] can be used. Note that Lines 10-11 in Algorithm 6 can be efficiently computed, and this is outlined in Algorithm 7.
Input : Two data tensors $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$ where $I_{1}\leq I_{2},\,I_{4}\leq I_{2}$ and a target tubal rank $R$
Output : A rank-$R$ GTCUR approximation for tensor pairs $(\underline{\bf X},\underline{\bf Y})$
$\underline{\bf X}$$\approx$ $\underline{\bf X}(:,{\bf p},:)*\underline{{\bf U}}_{1}*\underline{\bf X}({\bf s}_{1},:,:),$
$\quad\underline{\bf Y}$$\approx$ $\underline{\bf Y}(:,{\bf p},:)*\underline{{\bf U}}_{2}*\underline{\bf Y}({\bf s}_{2},:,:)$
1 $[\underline{\bf U},\underline{\bf V},\underline{\bf Z}]={\rm GTSVD}(\underline{\bf X},\underline{\bf Y})$;   (or randomized GTSVD [38])
2 for $j=1,2,\ldots,R$ do
3       ${\bf p}(j)={\arg\max}_{1\leq i\leq I_{2}}\|\underline{\bf Z}(i,j,:)\|$
4       ${\bf s}_{1}(j)=\arg\max_{1\leq i\leq I_{1}}\|\underline{\bf U}(i,j,:)\|$
5       ${\bf s}_{2}(j)=\arg\max_{1\leq i\leq I_{4}}\|\underline{\bf V}(i,j,:)\|$
6       $\underline{\bf Z}(:,j+1,:)=\underline{\bf Z}(:,j+1,:)-\underline{\bf Z}(:,1:j,:)*\underline{\bf Z}({\bf p},1:j,:)^{-1}*\underline{\bf Z}({\bf p},j+1,:)$
7       ${\underline{\bf U}(:,j+1,:)=\underline{\bf U}(:,j+1,:)-\underline{\bf U}(:,1:j,:)*\underline{\bf U}({\bf s}_{1},1:j,:)^{-1}*\underline{\bf U}({\bf s}_{1},j+1,:)}$
8       ${\underline{\bf V}(:,j+1,:)=\underline{\bf V}(:,j+1,:)-\underline{\bf V}(:,1:j,:)*\underline{\bf V}({\bf s}_{2},1:j,:)^{-1}*\underline{\bf V}({\bf s}_{2},j+1,:)}$
9
10 end for
[…]
${\underline{\bf U}}_{2}=\underline{\bf Y}(:,{\bf p},:)^{{\dagger}}*\underline{\bf Y}*\underline{\bf Y}({\bf s}_{2},:,:)^{{\dagger}}$
Algorithm 6 The GTCUR approximation for tensor pairs
Input : Two data tensors $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$ where $(I_{1}\leq I_{2},\,I_{4}\leq I_{2})$, and the index sets ${\bf p},\,{\bf s}_{1},\,{\bf s}_{2}$
Output : Computation of the middle tensors in Lines 10-11 of Algorithm 6 as
${\underline{\bf U}}^{(1)}=\underline{\bf X}(:,{\bf p},:)^{{\dagger}}*\underline{\bf X}*\underline{\bf X}({\bf s}_{1},:,:)^{{\dagger}}$
${\underline{\bf U}}^{(2)}=\underline{\bf Y}(:,{\bf p},:)^{{\dagger}}*\underline{\bf Y}*\underline{\bf Y}({\bf s}_{2},:,:)^{{\dagger}}$
1 $\widehat{\underline{\mathbf{X}}}={\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$;
2 $\widehat{\underline{\mathbf{Y}}}={\rm fft}\left({\underline{\mathbf{Y}},[],3}\right)$;
3 ${\underline{\mathbf{G}}}_{1}=\widehat{\underline{\mathbf{X}}}(:,{\bf p},:)$;
4 ${\underline{\mathbf{G}}}_{2}=\widehat{\underline{\mathbf{Y}}}(:,{\bf p},:)$;
5 ${\underline{\mathbf{G}}}_{3}=\widehat{\underline{\mathbf{X}}}({\bf s}_{1},:,:)$;
6 ${\underline{\mathbf{G}}}_{4}=\widehat{\underline{\mathbf{Y}}}({\bf s}_{2},:,:)$;
7 for $i=1,2,\ldots,\lceil\frac{I_{3}+1}{2}\rceil$ do
8       ${\underline{\mathbf{C}}}_{1}\left({:,:,i}\right)=\underline{\bf G}_{1}(:,:,i)\backslash(\widehat{{\bf X}}(:,:,i)/\underline{\bf G}_{3}({\bf s}_{1},:,i))$;
9       ${\underline{\mathbf{C}}}_{2}\left({:,:,i}\right)=\underline{\bf G}_{2}(:,:,i)\backslash(\widehat{{\bf X}}(:,:,i)/\underline{\bf G}_{4}({\bf s}_{2},:,i))$;
10
11 end for
12for $i=\lceil\frac{I_{3}+1}{2}\rceil+1,\ldots,I_{3}$ do
13       ${\underline{\mathbf{C}}}_{1}\left({:,:,i}\right)={\rm conj}({\underline{\mathbf{C}}}_{1}\left({:,:,I_{3}-i+2}\right))$;
14       ${\underline{\mathbf{C}}}_{2}\left({:,:,i}\right)={\rm conj}({\underline{\mathbf{C}}}_{2}\left({:,:,I_{3}-i+2}\right))$;
15 end for
[…]
$\underline{\mathbf{U}}^{(2)}={\rm ifft}\left({{\underline{\mathbf{C}}}_{2},[],3}\right)$;
Algorithm 7 Fast computations of Lines 10-11 in Algorithm 6
In what follows and similar to [2], we show the connection between the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ and TCUR of the tensors $\underline{\bf X}*\underline{\bf Y}^{-1}$ and $\underline{\bf X}*\underline{\bf Y}^{{\dagger}}$ for non-invertible tensor, or of full rank $\underline{\bf Y}$ in Theorem 3.
Theorem 3.
Let two data tensors $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$ be given, then
  • 1.
    a) If $\underline{\bf Y}$ is invertible, then the selected lateral and horizontal slice indices obtained based on the TCUR of the data tensor $\underline{\bf X}*\underline{\bf Y}^{-1}$ are identical to the index vectors ${\bf s}_{1}$ and ${\bf s}_{2}$ computed from the GTCUR of the tensor pair $(\underline{\bf X},\underline{\bf Y})$.
  • 2.
    b) As a special case, if $\underline{\bf Y}=\underline{\bf I}$, then the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ is the same as the TCUR of the data tensor $\underline{\bf X}$.
  • 3.
    c) If $\underline{\bf Y}$ is noninvertible but of full tubal rank, then the selected lateral and horizontal slice indices obtained via the TCUR of the data tensor $\underline{\bf X}*\underline{\bf Y}^{{\dagger}}$ are identical to the index vectors ${\bf s}_{1}$ and ${\bf s}_{2}$ computed via the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$.
Proof.
a) Let the GTSVD of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ be
$\underline{\bf X}$ $=$ $\underline{\bf U}*\underline{\bf C}*\underline{\bf Z}^{T},$ (16)
$\underline{\bf Y}$ $=$ $\underline{\bf V}*\underline{\bf S}*\underline{\bf Z}^{T},$ (17)
so, $\underline{\bf X}*\underline{\bf Y}^{-1}=\underline{\bf U}*(\underline{\bf C}*\underline{\bf S}^{-1})*\underline{\bf V}^{T}$ because $\underline{\bf Y}^{-1}=\underline{\bf Z}^{-T}*\underline{\bf S}^{-1}*\underline{\bf V}^{T}$. This means that the indices of the selected lateral and horizontal slices obtained via the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ are the same as those obtained from the TCUR of the tensor $\underline{\bf X}*\underline{\bf Y}^{-1}$ as both algorithms use the data tensors $\underline{\bf U}$ and $\underline{\bf V}$ to select the horizontal slices and lateral slices, respectively.
b) If $\underline{\bf Y}=\underline{\bf I}$, then from (17), we have $\underline{\bf Z}^{T}=\underline{\bf S}^{-1}*\underline{\bf V}^{T}$ and substituting it in (16), we arrive at $\underline{\bf X}=\underline{\bf U}*(\underline{\bf C}*\underline{\bf S}^{-1})*\underline{\bf V}^{T}$. Considering the same explanations mentioned for part a, the proof of this part is completed.
c) Since the data tensor $\underline{\bf Y}$ is of full tubal rank, its Moore-Penrose pseudoinverse can be represented as $\underline{\bf Y}^{{\dagger}}=\underline{\bf Z}^{-T}*\underline{\bf S}^{-1}*\underline{\bf V}^{T}$. So, the result is achieved using the same explanation as parts a and b. ∎
6 Tensor CUR approximation for tensor triplets based on the t-product
Figure 2: Visualization of the generalized tensor CUR (GTCUR) approximation for tensor triples $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$. The indices of the selected horizontal slices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$ are identical, but not those for the lateral slices. Also, the indices of the selected lateral slices of $\underline{\bf X}$ and $\underline{\bf Z}$ are the same, but not indices of the selected horizontal slices.
The MCUR approximation for matrix triplets was recently developed in [3] as a generalization of the MCUR. The idea is to use the restricted SVD (RSVD) [7] to select the indices of columns and rows. The RSVD decomposes a matrix relative to two other matrices, and indeed, acts on three matrices simultaneously. This idea can be generalized to tensors based on the t-SVD similarly and we call it GTCUR for tensor triplets. Here, we give the formal definition of this GTCUR approximation.
Definition 8.
Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, $\underline{\bf Y}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}}$, and $\underline{\bf Y}\in\mathbb{R}^{I_{5}\times I_{2}\times I_{3}}$ be given. The GTCUR approximation of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ with the tubal rank $R$ is defined as
$\underline{\bf X}$$\approx$ ${\underline{\bf C}}_{1}*\underline{\bf U}_{1}*\underline{\bf R}_{1}=(\underline{\bf X}*\underline{\bf P})*\underline{\bf U}_{1}*({\underline{\bf S}}^{T}*\underline{\bf X}),$ (18)
$\underline{\bf Y}$$\approx$ ${\underline{\bf C}}_{2}*\underline{\bf U}_{2}*\underline{\bf R}_{2}=(\underline{\bf Y}*\widehat{\underline{\bf P}})*\underline{\bf U}_{2}*({\underline{\bf S}}^{T}*\underline{\bf Y}),$ (19)
$\underline{\bf Z}$$\approx$ ${\underline{\bf C}}_{3}*\underline{\bf U}_{3}*\underline{\bf R}_{3}=(\underline{\bf Z}*\underline{\bf P})*\underline{\bf U}_{3}*(\widehat{\underline{\bf S}}^{T}*\underline{\bf Z}),$ (20)
where $\underline{\bf S}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\widehat{\underline{\bf S}}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf P}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ and $\widehat{\underline{\bf P}}\in\mathbb{R}^{I_{4}\times R\times I_{3}},\,R\leq\min(I_{1},I_{2},I_{4},I_{5})$ are the index selection tensors yielded by sampling some lateral/horizontal slices of the identity tensor according to the sampled lateral/horizontal slices, which are sampled from the data tensors $\underline{\bf X},\underline{\bf Y},$ and $\underline{\bf Z}$. Alternative representations for (18)-(20) are
$\underline{\bf X}$$\approx$ $\underline{\bf X}(:,{\bf p},:)*\underline{{\bf U}}_{1}*\underline{\bf X}({\bf s},:,:),$
$\underline{\bf Y}$$\approx$ $\underline{\bf Y}(:,{\bf p}_{1},:)*\underline{{\bf U}}_{2}*\underline{\bf Y}({\bf s},:,:),$
$\underline{\bf Z}$$\approx$ $\underline{\bf Z}(:,{\bf p},:)*\underline{{\bf U}}_{3}*\underline{\bf Z}({\bf s}_{1},:,:),$
where ${\bf p},\,{\bf s},\,{\bf s}_{1},\,{\bf p}_{1}$ are the index sets corresponding to the tensor selections $\underline{\bf S}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\widehat{\underline{\bf S}}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf P}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ and $\widehat{\underline{\bf P}}\in\mathbb{R}^{I_{4}\times R\times I_{3}}$.
The intuition behind the GTCUR for tensor triplets also comes from the tensor-based restricted t-SVD (t-RSVD) which is a direct generalization of the classical RSVD [7] to tensors and can be stated as follows. Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, $\underline{\bf Y}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}}$, and $\underline{\bf Z}\in\mathbb{R}^{I_{5}\times I_{2}\times I_{3}}$ be given. Generalizing the RSVD to tensors based on the t-product guarantees the following decompositions
$\underline{\bf X}=\underline{\bf L}*\underline{\bf D}_{1}*\underline{\bf W}^{T},\quad\underline{\bf Y}=\underline{\bf L}*\underline{\bf D}_{2}*\underline{\bf U}^{T},\quad\underline{\bf Z}=\underline{\bf V}*\underline{\bf D}_{3}*\underline{\bf W}^{T},$ (21)
where $\underline{\bf U}\in\mathbb{R}^{I_{2}\times I_{2}\times I_{3}},\,\underline{\bf V}\in\mathbb{R}^{I_{5}\times I_{5}\times I_{3}}$ are orthogonal and $\underline{\bf L}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}},\,\underline{\bf W}\in\mathbb{R}^{I_{2}\times I_{2}\times I_{3}}$ are nonsingular. Also, $\underline{\bf D}_{1}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}},\,\underline{\bf D}_{2}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}},\,\underline{\bf D}_{3}\in\mathbb{R}^{I_{5}\times I_{2}\times I_{3}}$ have quasi-diagonal frontal slices, which means that if we remove their zero columns and zero rows, they become diagonal. The tensor RSVD (t-RSVD) can be computed through the RSVD of the frontal slices of the underlying data tensors in the Fourier domain. To do so, we can replace the GSVD in Line 4 of Algorithm 5 with the RSVD. It is known that the computation of the RSVD of the frontal slices (matrices) can be performed using a double GSVD as described in [7, 39].
The procedure of the computation of the GTCUR for tensor triples is summarized in Algorithm 8. Lines 6-8 can be efficiently computed in the Fourier domain and similar algorithms like Algorithm 6 can be developed for this computation. The t-RSVD of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ provides the common tensor factors $\underline{\bf L},\,\underline{\bf W}$ and extra tensor factors $\underline{\bf U}$ and $\underline{\bf V}$. These bases can be used to sample lateral and horizontal slices of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$. Since we have the common tensor factors $\underline{\bf L}$ and $\underline{\bf W}$, the indices, which are used to sample the horizontal slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Z}$ are the same, while the indices for the selection of lateral slices of the data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ are identical. Nevertheless, the tensor bases $\underline{\bf U},\,\underline{\bf V}$ are used to sample lateral slice indices of the tensor $\underline{\bf Y}$ and horizontal slice indices of the tensor $\underline{\bf Z}$. The visualization of this approach is described in Figure 2. So, the idea is to compute the t-RSVD of the tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ and to find the corresponding tensor factors $\underline{\bf U},\,\underline{\bf V},\,\underline{\bf Z},\,\underline{\bf W}$ and use them to find the indices for selecting horizontal and lateral slices using the TDEIM algorithm. We should remark that the computation of the t-RSVD is demanding for the case of large-scale data tensors and the fast randomized GSVD algorithm proposed in [40] can be used for the computation of the double GSVD, which are required to compute the t-RSVD of the tensor triplets. The connection between the GTCUR for tensor triplets with the GTCUR for tensor pairs and the TCUR is described in the next theorem.
Theorem 4.
Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, $\underline{\bf Y}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}}$, and $\underline{\bf Z}\in\mathbb{R}^{I_{5}\times I_{2}\times I_{3}}$ be given, then
  • 1.
    a) If $\underline{\bf Y}$ and $\underline{\bf Z}$ are nonsingular tensors, then the selected lateral and horizontal slice indices from the TCUR decomposition of $\underline{\bf Y}^{-1}*\underline{\bf X}*\underline{\bf Z}^{-1}$ are identical to the index vectors ${\bf p}_{1}$ and ${\bf s}_{1}$, respectively, obtained from a GTCUR decomposition of the tensor triples $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$.
  • 2.
    b) In the particular case where $\underline{\bf Y}=\underline{{\bf I}}$ and $\underline{\bf Z}=\underline{{\bf I}}$, the GTCUR decomposition of $(\underline{\bf X},\underline{\bf I},\underline{\bf I})$ coincides with the TCUR decomposition of $\underline{\bf X}$.
  • 3.
    c) For a special choice of $\underline{\bf Y}=\underline{\bf I}$, a GTCUR approximation of $(\underline{\bf X},\underline{\bf I},\underline{\bf Z})$ coincides with the GTCUR decomposition of $(\underline{\bf X},\underline{\bf Z})$. Similar results can be stated for $\underline{\bf Z}=\underline{\bf I}$.
Proof.
a) From 21, it is not difficult to see that $\underline{\bf Y}^{-1}=\underline{\bf U}*{\underline{\bf D}}_{2}^{-1}*\underline{\bf L}^{-1},\,\underline{\bf Z}^{-1}=\underline{\bf W}*\underline{\bf D}_{3}^{-1}*\underline{\bf V}^{T}$, so we have
$\underline{\bf Y}^{-1}*\underline{\bf X}*\underline{\bf Z}^{-1}=(\underline{\bf U}*{\underline{\bf D}}_{2}^{-1}*\underline{\bf L}^{-1})*(\underline{\bf L}*\underline{\bf D}_{1}*\underline{\bf W}^{T})*(\underline{\bf W}*\underline{\bf D}_{3}^{-1}*\underline{\bf V}^{T})=$
$\underline{\bf U}*(\underline{\bf D}_{2}^{-1}\underline{\bf D}_{1}\underline{\bf D}_{3}^{-1})*\underline{\bf V}^{T}.$ (22)
This means that the selected selected lateral and horizontal slice indices of $\underline{\bf Y}^{-1}*\underline{\bf X}*\underline{\bf Z}^{-1}$ using the TCUR decomposition are the same as the index vectors ${\bf p}_{1}$ and ${\bf s}_{1}$, respectively, obtained from a GTCUR approximation of $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z}).$
b) If $\underline{\bf Y}=\underline{{\bf I}}$ and $\underline{\bf Z}=\underline{{\bf I}}$, then $\underline{\bf L}=\underline{\bf U}*{\underline{\bf D}}_{2}^{-1}$ and $\underline{\bf W}^{T}=\underline{\bf D}_{3}^{-1}*\underline{\bf V}^{T}$. Substituting them in the first part of (21), we have $\underline{\bf X}=\underline{\bf U}*(\underline{\bf D}_{2}^{-1}*\underline{\bf D}_{1}*\underline{\bf D}_{3}^{-1})*\underline{\bf V}^{T}$. This shows that both algorithms use the same tensors $\underline{\bf U}$ and $\underline{\bf V}$ to select the horizontal and lateral slices, respectively. This concludes the proof.
c) If $\underline{\bf Y}=\underline{{\bf I}}$, then from the second part of (21), we have $\underline{\bf L}=\underline{\bf U}*\underline{\bf D}_{2}^{-1}$. If we substitute this in the first and second parts of (21), we arrive at
$\underline{\bf X}$ $=$ $\underline{\bf U}*(\underline{\bf D}_{2}^{-1}*\underline{\bf D}_{1})*\underline{\bf W}^{T},$ (23)
$\underline{\bf Z}$ $=$ $\underline{\bf V}*\underline{\bf D}_{3}*\underline{\bf W}^{T}.$ (24)
Clearly, (23) and 24 shows that the GTCUR of the tensor pairs $(\underline{\bf X},\underline{\bf Z})$ provides the same indices for horizontal and lateral slice sampling as the GTCUR decomposition of the tensor pairs $(\underline{\bf X},\underline{\bf I},\underline{\bf Z})$. So, the proof is completed. ∎
Input : Three data tensors $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}},\,\underline{\bf Y}\in\mathbb{R}^{I_{1}\times I_{4}\times I_{3}},$ and $\underline{\bf Z}\in\mathbb{R}^{I_{5}\times I_{2}\times I_{3}}$; and a target tubal rank $R$
Output : A rank-$R$ GTCUR approximation of tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ as
$\underline{\bf X}$$\approx$ $\underline{\bf X}(:,{\bf p},:)*\underline{{\bf U}}_{1}*\underline{\bf X}({\bf s},:,:),$
$\underline{\bf Y}$$\approx$ $\underline{\bf Y}(:,{\bf p}_{1},:)*\underline{{\bf U}}_{2}*\underline{\bf Y}({\bf s},:,:)$
$\underline{\bf Z}$$\approx$ $\underline{\bf Z}(:,{\bf p},:)*\underline{{\bf U}}_{3}*\underline{\bf Z}({\bf s}_{1},:,:)$
1 Compute t-RSVD (either deterministic or randomized version) of $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ to obtain $\underline{\bf W},\,\underline{\bf Z},\underline{\bf U},\underline{\bf V}$;
2 ${\bf p}\leftarrow$ Apply Algorithm 4 to $\underline{\bf W}$;
3 ${\bf s}\leftarrow$ Apply Algorithm 4 to $\underline{\bf Z}$;
4 ${\bf p}_{1}\leftarrow$ Apply Algorithm 4 to $\underline{\bf U}$;
5 ${\bf s}_{1}\leftarrow$ Apply Algorithm 4 to $\underline{\bf V}$;
6 ${\underline{\bf U}}_{1}=\underline{\bf X}(:,{\bf p},:)^{{\dagger}}*\underline{\bf X}*\underline{\bf X}({\bf s},:,:)^{{\dagger}}$;
7 ${\underline{\bf U}}_{2}=\underline{\bf Y}(:,{\bf p}_{1},:)^{{\dagger}}*\underline{\bf Y}*\underline{\bf Y}({\bf s},:,:)^{{\dagger}}$;
8 ${\underline{\bf U}}_{3}=\underline{\bf Z}(:,{\bf p},:)^{{\dagger}}*\underline{\bf Z}*\underline{\bf Z}({\bf s}_{1},:,:)^{{\dagger}}$;
Algorithm 8 GTCUR approximation of tensor triplets
7 Conclusion
In this note, we showed how the tensor CUR (TCUR) approximation can be extended to tensor pairs and tensor triplets. We use the tensor Discrete Interpolatory Empirical Method (TDEIM) to generalize the TCUR to tensor pairs and tensor triplets. This extension is called the generalized TCUR (GTCUR) method. We established connections between some special cases of the GTCUR and the classical TCUR approximation. Efficient algorithms are presented to compute the GTCUR approximation for both tensor pairs and tensor triplets. We are investigating the theoretical and numerical aspects of the presented GTCUR approaches for tensor pairs and tensor triplets in a practical setting.
8 Conflict of Interest Statement
The author declares that he has no conflict of interest with anything.